# Energy-based generator matching: A neural sampler for general state space

### v2

@@ Appendix B Generator estimation in the multimodal spaces
Our estimator can also be applied to the mixed state spaces $S=X\times Y$ within the generator matching framework. Let $\{\tilde{p}_{t|1}(\cdot|x_{1})\}_{0\leq t\leq 1}$ and $\{\bar{p}_{t|1}(\cdot|y_{1})\}_{0\leq t\leq 1}$ denote the conditional probability paths on the $X$ and $Y$, respectively, and let ~$\tilde{\mathcal{L}}_{t|1}^{x_{1}}$ and ¯$\bar{\mathcal{L}}_{t|1}^{y_{1}}$ denote the corresponding conditional generators for $x_{1}\in X$ and $y_{1}\in Y$. Assume these generators are parameterized by $\tilde{F}_{t|1}^{x_{1}}:[0,1]\times S\to V_{1}$ and $\bar{F}_{t|1}^{y_{1}}:[0,1]\times S\to V_{2}$, respectively. For the joint space $S=X\times Y$, we consider the factorized conditional path:
@@
According to Proposition 5 in Holderrieth et al. [19], the conditional generator associated with $p_{t|1}$ admits the following parameterization:
@@ C.1 Generator of flow and jump model
In this section, we provide the definition of flow and discrete jump models, their generators and parametrizations. For the case of diffusion processes or more rigorous derivations, we refer the reader to Holderrieth et al. [19]. The discrete jump model is often referred to as a continuous-time Markov chain (CTMC).
@@
We derive the generator informally; see Davis [10] for a formal treatment:
@@
Remark on linear parametrization. Under mild regularity conditions (e.g., Feller processes), Holderrieth et al. [19] shows that Markov processes on both discrete and continuous state spaces can be universally expressed via linear parameterizations:
@@ C.2 Application to the conditional OT flow model
Using the identity from Equation 51, the estimated vector field $u_{t}(x_{t})$ becomes:
@@
where $x_{1}^{(i)}\sim q_{1|t}(\cdot|x_{t})$. This is precisely the same estimator used in Woo and Ahn [39] for the flow-based sampler.
Assumption check for bootstrapping. Next, we derive the bootstrapping estimator. We construct the backward transition kernel $p_{t|r}$ satisfying the marginal consistency Equation 52:
@@
[…]
@@
Now, verify that the conditional vector field $u_{t|r}$ satisfies the marginal consistency Equation 53:
@@
which implies the proposed transition kernel $p_{t|r}(x_{t}|x_{r})$ and conditional vector field $u_{t|r}(x_{t}|x_{r})$ satisfies the assumption of our Theorem 1.
@@ C.3 Application to the masked diffusion model
This yields the simple importance weight $\tilde{w}(x_{t},x_{1})=\tilde{p}_{1}(x_{1})/Z_{1|t}(x_{t})$. Following Equation 51, the estimator for the transition matrix $u_{t}(y_{t},x_{t})$ becomes:
@@
In practice, the state space $S=[N]^{D}$ factorizes along dimensions, where $D$ is sequence length and $[N]=\{1,\dots,N\}$. We thus factorize the masked diffusion path as follows:
@@
Assumption check for bootstrapping. The backward transition kernel $p_{t|r}$ of masked diffusion satisfies the marginal consistency since it defines the Markov process (noising process of masked diffusion). Thus, it is suffice to show that the conditional transition rate matrix $u_{t|r}(y_{t},x_{t}|x_{r})$ satisfies the marginal consistency Equation 53:
@@
[…]
@@
=$=\begin{cases}0&(x_{t}=x_{1})\\ \frac{\dot{\kappa_{t}}}{(\kappa_{r}-\kappa_{t})(1-\kappa_{t})}\left(\delta_{x_{1}}(y_{t})-\delta_{x_{t}}(y_{t})\right)\left(1-\frac{\kappa_{t}}{\kappa_{r}}\right)\kappa_{r}&(x_{t}=M)\end{cases}$ (165)
@@
where we used the fact that $p_{r|1}(x_{r}|x_{1})>0$ for only $x_{r}=M$ or $x_{r}=x_{1}$ in the fourth equality. Hence, the proposed transition kernel $p_{t|r}$ and conditional transition rate matrix $u_{t|r}$ satisfies the assumption of our Theorem 1.
@@
=$=\begin{cases}\frac{\kappa_{t}}{\kappa_{r}}&(x_{t}\neq M,x_{r}=x_{t})\\ 0&(x_{t}\neq M,x_{r}\neq x_{t})\\ 1-\frac{\kappa_{t}}{\kappa_{r}}&(x_{t}=M,x_{r}\neq M)\\ 1&(x_{t}=M,x_{r}=M)\end{cases}$ (169)
@@
$q_{r|t}(x_{r}|x_{t})=\begin{cases}\delta_{x_{t}}(x_{r})&(x_{t}\neq M)\\ \text{Cat}(1-\frac{\kappa_{t}}{\kappa_{r}},\dots,1-\frac{\kappa_{t}}{\kappa_{r}},1)&(x_{t}=M)\\ \end{cases}$ (170)
@@
=$=\begin{cases}(N-1)(1-\kappa_{t})&(x_{r}=M)\\ \kappa_{r}&(x_{r}\neq M)\end{cases}$ (175)
@@ D.1 Task details
Discrete Ising model. We consider the Ising model defined on a 2D grid $\{-1,1\}^{L\times L}$ with size $L$. The energy function :$\mathcal{E}:\{-1,1\}^{L\times L}\to\mathbb{R}$ is given by:
@@
where $x_{1},a\in\mathbb{R}^{2}$, $x_{2},b\in\{0,1\}^{3}$, $\Sigma\in\mathbb{R}$, and $W\in\mathbb{R}^{2\times 3}$. Parameters are selected to induce multiple modes, specifically six modes in continuous dimensions (see Figure 5). We set:
@@ D.2 Metrics
Evaluation metrics in our experiments primarily utilize Wasserstein distances, computed via the Python Optimal Transport (POT) library [13] using exact linear programming. Specifically, we measure the distances between 2000 empirical samples generated by our samplers and 2000 ground truth samples uniformly selected from extensive Gibbs sampling or exact sampling processes.
@@ D.3 Experimental setup
Ising $0.25\times 5,\ \beta=$ 0.2 EGM 256 3 - -
@@
Ising $0.45\times 5,\ \beta=$ 0.4 EGM 256 3 - -
@@
Ising $0.210\times 10,\ \beta=$ 0.2 EGM 1024 3 - -
@@
Ising $0.410\times 10,\ \beta=$ 0.4 EGM 256 3 - -
@@
As a baseline, we report the performance of a traditional Gibbs sampler [16]. Specifically, we ran Gibbs sampling with four parallel chains, each performing 6000 steps, collecting a total of 24,000 samples. For evaluation purposes, we uniformly subsampled 2000 samples from this set.
@@
Additional task-specific details are provided below, and optimal hyperparameters are summarized in Table 4.
Discrete Ising model. We employed a 3-layer MLP with sinusoidal time embeddings for both the intermediate energy function and the transition rate matrix. Each discrete token representing spin values -1 or 1 was embedded in 4 dimensions. Following Gat et al. [14], the transition rate matrix $u_{t}(y,x)$ was parametrized using a probability denoiser $p_{1|t}(y|x)$ analogous to the $x_{1}$-prediction in the flow models. Hidden dimensions were explored within $\{256,512\}$, with additional trials at $\{1024,2048\}$ for the $10\times 10$ Ising grid.
GB-RBM. We utilized a 6-layer residual MLP with 128 hidden units, a 4-dimensional discrete embedding, and a 64-dimensional continuous embedding. Discrete and continuous embeddings were concatenated and fed into the shared 6-layer MLP. Separate predictor networks subsequently estimated the continuous drift and discrete transition rate matrix. The conditional OT path performed best for both EGM and bootstrapping. We clipped the regression target $F_{t}$ at a maximum norm of 20 and the energy estimator ^$\hat{\mathcal{E}}_{t}$ at 100 to stabilize training.
JointDW4. The network architecture matched that used in GB-RBM. The conditional OT path again yielded optimal performance for both methods. Regression targets $F_{t}$ and energy estimates ^$\hat{\mathcal{E}}_{t}$ were clipped at maximum norms of 100 and 1000, respectively.
@@ Appendix E Limitations and Discussion
Comparison to LEAPS. We compare our method to LEAPS [20], a neural sampler designed for discrete spaces. Our framework is more general in that it applies to arbitrary state spaces and Markov processes, including both continuous and discrete cases, whereas LEAPS is limited to discrete domains. Even when instantiated with a discrete sampler, EGM and LEAPS differ fundamentally. EGM relies on the prescribed conditional probability paths that mix the target distribution, while LEAPS is built on the escorted transport with a temperature annealing. In continuous domains, it has been shown that geometric annealing paths can lead to optimal drifts with high Lipschitz constants [27], which limits sampler performance; whether a similar issue arises in discrete spaces remains an open question.
@@ F.1 Additional qualitative results
We provide additional qualitative results for the experiments in Figure 2. In Figure 6, we plot the energy histogram of the GB-RBM and JointDW4 compared to the ground truth sample. The Gibbs sampler baseline, EGM, and Bootstrapping match the ground truth energy histogram. However, the Gibbs sampler on GB-RBM suffers from mode collapse as demonstrated in Figure 3.
@@ F.2 Effective sample size of our MC estimator
where $\tilde{w}_{i}$ denotes the unnormalized importance weight with the $i$-th proposed sample and $n$ is the total number of MC samples. We report the normalized ESS to indicate the fraction of effectively used samples. Figure 7 shows the average normalized ESS over the course of training. The bootstrapped estimator maintains a significantly higher ESS during training, confirming its improved utilization of proposed samples compared to the naive EGM.

### v3

@@ Appendix B Generator estimation in the multimodal spaces
Our estimator can also be applied to the mixed state spaces $S=X\times Y$ within the generator matching framework. Let $\{\tilde{p}_{t|1}(\cdot|x_{1})\}_{0\leq t\leq 1}$ and $\{\bar{p}_{t|1}(\cdot|y_{1})\}_{0\leq t\leq 1}$ denote the conditional probability paths on the $X$ and $Y$, respectively, and let $\tilde{\mathcal{L}}_{t|1}^{x_{1}}$ and $\bar{\mathcal{L}}_{t|1}^{y_{1}}$ denote the corresponding conditional generators for $x_{1}\in X$ and $y_{1}\in Y$. Assume these generators are parameterized by $\tilde{F}_{t|1}^{x_{1}}:[0,1]\times S\to V_{1}$ and $\bar{F}_{t|1}^{y_{1}}:[0,1]\times S\to V_{2}$, respectively. For the joint space $S=X\times Y$, we consider the factorized conditional path:
@@
According to Proposition 5 in holderrieth2024generator, the conditional generator associated with $p_{t|1}$ admits the following parameterization:
@@ C.1 Generator of flow and jump model
In this section, we provide the definition of flow and discrete jump models, their generators and parametrizations. For the case of diffusion processes or more rigorous derivations, we refer the reader to holderrieth2024generator. The discrete jump model is often referred to as a continuous-time Markov chain (CTMC).
@@
We derive the generator informally; see davis1984piecewise for a formal treatment:
@@
Remark on linear parametrization. Under mild regularity conditions (e.g., Feller processes), holderrieth2024generator shows that Markov processes on both discrete and continuous state spaces can be universally expressed via linear parameterizations:
@@ C.2 Application to the conditional OT flow model
Using the identity from Equation˜51, the estimated vector field $u_{t}(x_{t})$ becomes:
@@
where $x_{1}^{(i)}\sim q_{1|t}(\cdot|x_{t})$. This is precisely the same estimator used in woo2024iterated for the flow-based sampler.
Assumption check for bootstrapping. Next, we derive the bootstrapping estimator. We construct the backward transition kernel $p_{t|r}$ satisfying the marginal consistency Equation˜52:
@@
$\overset{d}{=}tX_{1}+(1-t)\epsilon^{\prime}_{t},\quad\epsilon^{\prime}_{t}\sim\mathcal{N}(0,I),$ (145)
where $\overset{d}{=}$ denotes that two random variables have same distribution. Thus, marginal consistency holds.
@@
Now, verify that the conditional vector field $u_{t|r}$ satisfies the marginal consistency Equation˜53:
@@
which implies the proposed transition kernel $p_{t|r}(x_{t}|x_{r})$ and conditional vector field $u_{t|r}(x_{t}|x_{r})$ satisfies the assumption of our Theorem˜1.
@@ C.3 Application to the masked diffusion model
This yields the simple importance weight $\tilde{w}(x_{t},x_{1})=\tilde{p}_{1}(x_{1})/Z_{1|t}(x_{t})$. Following Equation˜51, the estimator for the transition matrix $u_{t}(y_{t},x_{t})$ becomes:
@@
In practice, the state space $S=[N]^{D}$ factorizes along dimensions, where $D$ is sequence length and $N]=\{1,\dots,N\}$. We thus factorize the masked diffusion path as follows:
@@
Assumption check for bootstrapping. The backward transition kernel $p_{t|r}$ of masked diffusion satisfies the marginal consistency since it defines the Markov process (noising process of masked diffusion). Thus, it is suffice to show that the conditional transition rate matrix $u_{t|r}(y_{t},x_{t}|x_{r})$ satisfies the marginal consistency Equation˜53:
@@
$(\text{L.H.S.})$ $=\sum_{x_{r}}\frac{\dot{\kappa_{t}}}{\kappa_{r}-\kappa_{t}}\left(\delta_{x_{r}}(y_{t})-\delta_{x_{t}}(y_{t})\right)p_{r|1,t}(x_{r}|x_{1},x_{t})$ (187)
@@
$=\begin{cases}0&(x_{t}=x_{1})\\ \frac{\dot{\kappa_{t}}}{(\kappa_{r}-\kappa_{t})(1-\kappa_{t})}\left(\delta_{x_{1}}(y_{t})-\delta_{x_{t}}(y_{t})\right)\left(1-\frac{\kappa_{t}}{\kappa_{r}}\right)\kappa_{r}&(x_{t}=M)\end{cases}$ (194)
@@
where we used the fact that $p_{r|1}(x_{r}|x_{1})>0$ for only $x_{r}=M$ or $x_{r}=x_{1}$ in the fourth equality. Hence, the proposed transition kernel $p_{t|r}$ and conditional transition rate matrix $u_{t|r}$ satisfies the assumption of our Theorem˜1.
@@
$=\begin{cases}\frac{\kappa_{t}}{\kappa_{r}}&(x_{t}\neq M,x_{r}=x_{t})\\ 0&(x_{t}\neq M,x_{r}\neq x_{t})\\ 1-\frac{\kappa_{t}}{\kappa_{r}}&(x_{t}=M,x_{r}\neq M)\\ 1&(x_{t}=M,x_{r}=M)\end{cases}$ (198)
@@
[…]
@@
$=\begin{cases}(N-1)(1-\kappa_{t})&(x_{r}=M)\\ \kappa_{r}&(x_{r}\neq M)\end{cases}$ (204)
@@ D.1 Task details
Discrete Ising model. We consider the Ising model defined on a 2D grid $\{-1,1\}^{L\times L}$ with size $L$. The energy function $\mathcal{E}:\{-1,1\}^{L\times L}\to\mathbb{R}$ is given by:
@@
where $x_{1},a\in\mathbb{R}^{2}$, $x_{2},b\in\{0,1\}^{3}$, $\Sigma\in\mathbb{R}$, and $W\in\mathbb{R}^{2\times 3}$. Parameters are selected to induce multiple modes, specifically six modes in continuous dimensions (see Figure˜5). We set:
@@ D.2 Metrics
Evaluation metrics in our experiments primarily utilize Wasserstein distances, computed via the Python Optimal Transport (POT) library [flamary2021pot] using exact linear programming. Specifically, we measure the distances between 2000 empirical samples generated by our samplers and 2000 ground truth samples uniformly selected from extensive Gibbs sampling or exact sampling processes.
@@ D.3 Experimental setup
Ising $5\times 5,\ \beta=0.2$ EGM 256 3 - -
@@
Ising $5\times 5,\ \beta=0.4$ EGM 256 3 - -
@@
Ising $10\times 10,\ \beta=0.2$ EGM 1024 3 - -
@@
Ising $10\times 10,\ \beta=0.4$ EGM 256 3 - -
@@
As a baseline, we report the performance of a traditional Gibbs sampler [geman1984stochastic]. Specifically, we ran Gibbs sampling with four parallel chains, each performing 6000 steps, collecting a total of 24,000 samples. For evaluation purposes, we uniformly subsampled 2000 samples from this set.
@@
Additional task-specific details are provided below, and optimal hyperparameters are summarized in Table˜4.
Discrete Ising model. We employed a 3-layer MLP with sinusoidal time embeddings for both the intermediate energy function and the transition rate matrix. Each discrete token representing spin values -1 or 1 was embedded in 4 dimensions. Following gat2024discrete, the transition rate matrix $u_{t}(y,x)$ was parametrized using a probability denoiser $p_{1|t}(y|x)$ analogous to the $x_{1}$-prediction in the flow models. Hidden dimensions were explored within $\{256,512\}$, with additional trials at $\{1024,2048\}$ for the $10\times 10$ Ising grid.
GB-RBM. We utilized a 6-layer residual MLP with 128 hidden units, a 4-dimensional discrete embedding, and a 64-dimensional continuous embedding. Discrete and continuous embeddings were concatenated and fed into the shared 6-layer MLP. Separate predictor networks subsequently estimated the continuous drift and discrete transition rate matrix. The conditional OT path performed best for both EGM and bootstrapping. We clipped the regression target $F_{t}$ at a maximum norm of 20 and the energy estimator $\hat{\mathcal{E}}_{t}$ at 100 to stabilize training.
JointDW4. The network architecture matched that used in GB-RBM. The conditional OT path again yielded optimal performance for both methods. Regression targets $F_{t}$ and energy estimates $\hat{\mathcal{E}}_{t}$ were clipped at maximum norms of 100 and 1000, respectively.
@@ Appendix E Limitations and Discussion
Comparison to LEAPS. We compare our method to LEAPS [holderrieth2025leaps], a neural sampler designed for discrete spaces. Our framework is more general in that it applies to arbitrary state spaces and Markov processes, including both continuous and discrete cases, whereas LEAPS is limited to discrete domains. Even when instantiated with a discrete sampler, EGM and LEAPS differ fundamentally. EGM relies on the prescribed conditional probability paths that mix the target distribution, while LEAPS is built on the escorted transport with a temperature annealing. In continuous domains, it has been shown that geometric annealing paths can lead to optimal drifts with high Lipschitz constants [matelearning], which limits sampler performance; whether a similar issue arises in discrete spaces remains an open question.
@@ F.1 Additional qualitative results
We provide additional qualitative results for the experiments in Section˜4. In Figure˜6, we plot the energy histogram of the GB-RBM and JointDW4 compared to the ground truth sample. The Gibbs sampler baseline, EGM, and Bootstrapping match the ground truth energy histogram. However, the Gibbs sampler on GB-RBM suffers from mode collapse as demonstrated in Figure˜3.
@@ F.2 Effective sample size of our MC estimator
where $\tilde{w}_{i}$ denotes the unnormalized importance weight with the $i$-th proposed sample and $n$ is the total number of MC samples. We report the normalized ESS to indicate the fraction of effectively used samples. Figure˜7 shows the average normalized ESS over the course of training. The bootstrapped estimator maintains a significantly higher ESS during training, confirming its improved utilization of proposed samples compared to the naive EGM.
@@
F.3 Additional quantitative results
Comparison to additional baselines.
Because our proposed method is applicable to any state space and any Markov process, it can also be directly applied to purely continuous settings, where most neural samplers have been developed. Although many of these existing methods are specialized for continuous domains, we include them as references to contextualize our results. For discrete settings, we additionally include specialized baselines to further demonstrate the performance of our approach. As emphasized earlier, prior neural samplers are typically designed for either continuous or discrete domains and often rely on domain-specific architectures to achieve efficiency, whereas our method imposes no such constraints.
For continuous tasks, we compare against iDEM and PIS on the ManyWell32 benchmark [sendera2024improved]. For discrete tasks, we compare against LEAPS on a $15\times 15$ Ising model.
Table 5: Performance of EGM compared to PIS and iDEM on ManyWell32. BS denotes bootstrapping.
Algorithm -$\mathcal{E}\text{-}\mathcal{W}_{1}$ $\downarrow$ $x\text{-}\mathcal{W}_{2}$ $\downarrow$ # of modes
EGM + BS 3.30 7.07 4
PIS 3.39 6.14 1
iDEM 5.53 7.74 4
Table 6: Performance of EGM compared to LEAPS on the Ising model ($15\times 15$, $\beta=0.28$). BS denotes bootstrapping.
Algorithm -$\mathcal{E}\text{-}\mathcal{W}_{1}$ $\downarrow$ $M\text{-}\mathcal{W}_{1}$ $\downarrow$
EGM 0.89 0.06
EGM + BS 0.65 0.04
LEAPS 0.68 0.01
LEAPS + MCMC 0.49 0.01
The results show that EGM matches or slightly outperforms existing diffusion-based samplers (iDEM, PIS) in continuous settings. In the discrete case, LEAPS performs comparably to EGM+BS (bootstrapping). Notably, LEAPS combined with MCMC achieves slightly better performance than EGM+BS, albeit at the cost of significantly more energy evaluations (45B vs. 85B).
Computational complexity analysis.
For reference, Table 7 summarizes the computational complexity of EGM compared to a classical Gibbs sampler. We report the number of energy evaluations, wall-clock time per training epoch, and GPU memory usage on a $10\times 10$ Ising model with $\beta=0.4$. All experiments are conducted on an NVIDIA RTX 3090 GPU.
Table 7: Computational complexity of EGM compared to the Gibbs sampler.
Method # Energy evals Wall-clock time (1 epoch) Memory footprint
EGM 200M 0.16 s 12 GB
EGM + BS 200M 0.038 s 12 GB
Parallel Gibbs 280K 1 GB
Effect of bootstrapping time step.
We also study how performance varies with the time gap $\epsilon=r-t>0$ between two time steps $r$ and $t$. Table 8 reports the results on a $5\times 5$ Ising model with $\beta=0.4$.
Table 8: Effect of the bootstrapping time gap $\epsilon$ on performance.
$\epsilon$ -$\mathcal{E}\text{-}\mathcal{W}_{1}$ $\downarrow$ $M\text{-}\mathcal{W}_{1}$ $\downarrow$
0.01 2.84 0.32
0.02 0.96 0.09
0.05 0.84 0.08
0.10 0.58 0.08
0.20 1.90 0.09
0.50 3.73 0.20
As $\epsilon$ decreases, bootstrapping generally improves performance. However, if $\epsilon$ becomes too small, the scale of the conditional generator $F_{t|r}$ can grow rapidly, causing large fluctuations in loss magnitude across time steps. This leads to unstable neural network optimization due to inconsistent gradient scales. Consequently, choosing a moderately small $\epsilon$ is critical for stable training.
EMA and training stability.
Finally, we examine the effect of applying an exponential moving average (EMA) to the parameters of the energy model. EMA is a standard stabilization technique in reinforcement learning, used to mitigate the moving-target problem when regressing on a learned value function. A similar effect is observed here: applying EMA significantly stabilizes training when the generator is conditioned on a learned energy model. Table 9 shows results with and without EMA on a $10\times 10$ Ising model with $\beta=0.4$. All other hyperparameters are fixed.
Table 9: Effect of EMA on training stability.
-$\mathcal{E}\text{-}\mathcal{W}_{1}$ $\downarrow$ $M\text{-}\mathcal{W}_{1}$ $\downarrow$
[…]
Using EMA substantially reduces the variance of both energy-based and magnetization metrics, confirming its effectiveness in improving training stability.